OpenStudy (anonymous):

dy/dx = y/(x+y^2cos y)

OpenStudy (jamesj):

This problem really belongs in mathematics, but it's so good I'll try and answer it here. It's easiest to think of this as dx/dy = (x + y^2 cos y)/y = x/y + y cos y or x' - (1/y)x = y cos y Given that form, you can solve using standard first-order ODE techniques.

OpenStudy (jamesj):

i.e., x(y) = Ay + By.sin y Now you need to invert that equation. As you can see, the inverse in general won't be pretty.

OpenStudy (anonymous):

thanks dude

OpenStudy (anonymous):

invert the already inverted one?

OpenStudy (jamesj):

I mean you'd need to find the inverse function of x(y).

OpenStudy (anonymous):

please explain

OpenStudy (jamesj):

Your original equation implicitly asks for a solution y = f(x), y as a function of x. The solution I've provided gives x as a function of y. Hence if you actually need y = f(x), then you'll need to find the inverse function of x = g(y) = Ay + By.sin y The inverse function of g will be the function f. \[ f^{-1} = g \]

OpenStudy (anonymous):

Thanks again homie